OpenStudy (anonymous):

Find the derivative... (t)/((t^2+t^9)(t^5+1))

OpenStudy (caozeyuan):

what a mess!

OpenStudy (caozeyuan):

do you mean something like\[\frac{ t }{ (t ^{2}+t ^{9})\times(t ^{5} +1) }\]

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

you have to use quotient rule

OpenStudy (caozeyuan):

\[\left( \frac{ u }{ v } \right)^{\prime}= \frac{ u ^{\prime}v-v ^{\prime}u }{ v^{2} }\]

OpenStudy (caozeyuan):

here u=t and v is the huge stuff at bottom